how do u write 3 equivalent ratios for 2/9
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

write 3 equivalent ratios for 2/9
-------
(2/9)(2/2) = 4/18
---
(2/9)(3/3) = 6/27
---
(2/9)(4/4) = 8/36
======================
Cheers,
Stan H.

Answer by richwmiller(17219)

4/18 by multiplying by 2/2
6/27 by multiplying by 3/3
8/36 by multiplying by 4/4
10/45 by multiplying by 5/5
12/54 by multiplying by 6/6